How this could be SLOVE??
sqrt( 4x^2 +5-1 )=2(x+2)

Answer by checkley77(12844)

sqrt(4x^2+5-1)=2(x+2)
sqrt(4x^2+4)=2x+4
now square both sides:
4x^2+4=(2x+4)^2
4x^2+4=4x^2+16x+16
4x^2-4x^2-16x=16-4
-16x=12
x=12/-16
x=-3/4 answer.